OpenStudy (anonymous):

Simplify sqrt(18(s^12)(t^5))

OpenStudy (anonymous):

how many times does 2 go in to 12?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

s^6+6?

OpenStudy (anonymous):

\[\sqrt{16s ^{12}t ^{5}}\]

OpenStudy (anonymous):

is what the equation looks like

OpenStudy (anonymous):

oops 18**

OpenStudy (anonymous):

sqrt(18(s^12)(t^5))=(18(s^12)(t^5))^(1/2) =18^(1/2)*s^(6)*t^(5/2) =3*sqrt(2t^5)*s^6

OpenStudy (anonymous):

thx